Q1]

A corporation's stock promises to pay a dividend of $5 next year and investors expect the price of this stock next year to be $31. Because of the risk associated with this stock, investors require a rate of return of 20 percent from it. The price of this stock in the stock market is ( )dollars.

No comma, dollar sign, or decimal places.

Q2]


There are two assets in the market. They both promise to pay $1,120 next year. Asset A is less risky and carries a rate of return of 12 percent. The risk premium on Asset B is 16 percent. Therefore,

Price of Asset A = ( ) dollars

Price of Asset B =( ) dollars.

No commas, decimal places, or dollar signs.
1
Expert's answer
2017-05-03T09:04:08-0400
Q1] P=5/0.2
p=25
Q2] A)P=1120/0.12
P=9333
B) P=1120/0.16
P=7000